OpenStudy (anonymous):

How do you find the derivative of natural logs? Find the derivative h(z)=(ln9)^z Thanks.

OpenStudy (phi):

I assume you want to take the derivative with respect to z. use the idea \[ e^{\ln(a)}= a\] in this case, let a= ln(9) so \[ \ln(9) = e^{\ln( \ln(9))}\] rewrite your expression as \[ \left( e^{\ln( \ln(9))}\right)^z = e^{\ln( \ln(9)) z}\] now that might look ugly, but ln(ln(9)) is just an (ugly) constant number so you are taking the derivative of \[ \frac{ d\ e^{bz}}{dz} \] where b is a constant= ln(ln(9))

OpenStudy (anonymous):

Thanks, but I'm not getting the correct answer when I input e^ln(ln(9))z

OpenStudy (phi):

that's because the \[ \frac{d}{dz} e^{bz} = e^{bz} \frac{d}{dz}bz = b e^{bz} \]

OpenStudy (phi):

so the answer is \[ \ln(\ln(9)) e^{\ln(\ln(9))z} \]

OpenStudy (phi):

they may re-write e^(ln(ln(9)) as ln(9) so another way to write the answer is \[ \ln(\ln(9)) (\ln(9))^z \]

OpenStudy (anonymous):

Oohhh thank you! That makes much more sense now. But why isn't z taken down as well?